OpenStudy (anonymous):

The sum of two numbers is x. If one of the numbers is 12, what is the other? 12 - x x - 12 12 + x 12 · x

Parth (parthkohli):

12 - x of course!

Parth (parthkohli):

First option..

OpenStudy (zzr0ck3r):

x= 12 + y so subtract 12 from both sides to know what y is x-12=y

OpenStudy (anonymous):

a+b=x, where a and b are numbers if b=12 then a=x-12 works the same for way for a=12